OpenStudy (anonymous):

I'm doing 3.15 x 2.2 I keeping getting 69.30 the answer key says 6.93....how? what am I doing wrong?

OpenStudy (anonymous):

\[\frac{ 315 }{ 100 } \times \frac{ 22 }{ 10 }= \frac{ 6930 }{ 1000 }= 6.930\]

OpenStudy (anonymous):

You are plugging it into the calculator wrong. Try plugging in 3.15 first. Then hit the multiply button. Then hit 2.2 Then hit the equal button. Tell me what you get?

OpenStudy (mathstudent55):

3 times 2 is 6, so 3.15 times 2.22 can't be much more than 6. It cannot possibly be 69.30

OpenStudy (anonymous):

OMG I CAN'T BREATHE

OpenStudy (mathstudent55):